# Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis

### v2

@@ Abstract
$\Lambda^{+}_{c}\to p\pi^{0}$ $0.0156(75)$ $0.0174(53)$ $0.0164(64)$
$\Lambda^{+}_{c}\to pK_{S}^{0}$ $1.59(7)$ $0.2(5)$ $1.646(48)$ $0.41(12)$ $1.634(49)$ $0.12(26)$
$\Lambda^{+}_{c}\to pK_{L}^{0}$ $1.67(7)$ $1.646(48)$ $1.634(49)$
$\Lambda^{+}_{c}\to p\eta$ $0.158(11)$ $0.158(11)$ $0.155(11)$
$\Lambda^{+}_{c}\to p\eta^{\prime}$ $0.048(9)$ $0.0464(61)$ $0.0541(76)$
$\Lambda^{+}_{c}\to\Lambda\pi^{+}$ $1.29(5)$ $-0.755(6)$ $1.305(47)$ $-0.7538(60)$ $1.290(49)$ $-0.7550(60)$
$\Lambda^{+}_{c}\to\Sigma^{0}\pi^{+}$ $1.27(6)$ $-0.466(18)$ $1.259(46)$ $-0.471(15)$ $1.261(49)$ $-0.472(15)$
$\Lambda^{+}_{c}\to\Sigma^{+}\pi^{0}$ $1.24(9)$ $-0.484(27)$ $1.273(46)$ $-0.469(15)$ $1.272(49)$ $-0.471(15)$
$\Lambda^{+}_{c}\to\Xi^{0}K^{+}$ $0.55(7)$ $0.01(16)$ $0.417(29)$ $0.549(69)$ 0.02(15)
$\Lambda^{+}_{c}\to\Lambda^{0}K^{+}$ $0.0642(31)$ $-0.58(5)$ $0.0641(29)$ $-0.556(44)$ $0.0643(31)$ $-0.579(50)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta$ $0.32(5)$ $-0.99(6)$ $0.327(49)$ $-0.9998(35)$ $0.321(50)$ $-0.973(57)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.41(8)$ $-0.46(7)$ $0.416(65)$ $-0.455(64)$ $0.354(66)$ $-0.446(66)$
$\Lambda^{+}_{c}\to\Sigma^{0}K^{+}$ $0.0370(31)$ $-0.54(20)$ $0.0376(18)$ $-0.9971(36)$ $0.0363(29)$ $-0.54(17)$
$\Lambda^{+}_{c}\to n\pi^{+}$ $0.066(13)$ $0.0641(23)$ $0.063(11)$
$\Lambda^{+}_{c}\to\Sigma^{+}K_{S}^{0}$ $0.047(14)$ $0.0311(29)$ $0.043(11)$
$\Xi^{+}_{c}\to\Xi^{0}\pi^{+}$ $1.6(8)$ $0.880(78)$ $1.90(68)$
$\Xi^{0}_{c}\to\Lambda K_{S}^{0}$ $0.32(6)$ $0.248(29)$ $0.331(59)$
$\Xi^{0}_{c}\to\Xi^{-}\pi^{+}$ $1.43(27)$ $-0.640(51)$ $1.17(18)$ $-0.701(45)$ $1.25(20)$ $-0.640(51)$
$\Xi^{0}_{c}\to\Xi^{-}K^{+}$ $0.039(11)$ $0.0520(80)$ $0.0485(74)$
@@
$\Lambda^{+}_{c}\to\Sigma^{+}\eta$ $0.327(49)$ $-0.9998(35)$ $0.327(85)$ $-0.95(19)$ $-0.3(1.7)$ $0.1(3.3)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.416(65)$ $-0.455(64)$ $0.38(10)$ $-0.442(82)$ $-0.70(97)$ $0.6(1.2)$
$\Lambda^{+}_{c}\to p\eta$ $0.154(10)$ $0.885(78)$ $0.162(23)$ $0.35(83)$ $-0.07(18)$ $-0.94(31)$ $-0.047(45)$
$\Lambda^{+}_{c}\to p\eta^{\prime}$ $0.0466(61)$ $-0.990(16)$ $0.059(20)$ $-0.82(63)$ $-0.6(1.1)$ $0.12(71)$ $-0.003(275)$
$\Xi^{+}_{c}\to\Sigma^{+}\eta$ $0.116(14)$ $0.91(11)$ $0.09(12)$ $0.3(1.0)$ $-0.1(1.1)$ $-0.93(23)$ $-0.08(10)$
$\Xi^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.117(16)$ $-0.418(71)$ $0.137(67)$ $-0.20(20)$ $-0.87(91)$ $0.4(1.8)$ $0.03(16)$
$\Xi^{+}_{c}\to p\eta$ $0.00908(67)$ $-0.054(61)$ $0.0107(70)$ $-0.15(21)$ $0.48(39)$ $-0.86(25)$
$\Xi^{+}_{c}\to p\eta^{\prime}$ $0.0091(10)$ $-0.9957(86)$ $0.0096(92)$ $-0.83(61)$ $-0.56(80)$ $-0.03(1.79)$
$\Xi^{0}_{c}\to\Xi^{0}\eta$ $0.177(75)$ $1.000(98)$ $0.082(68)$ $0.8(1.6)$ $0.5(9.8)$ $-0.3(14.9)$
$\Xi^{0}_{c}\to\Xi^{0}\eta^{\prime}$ $0.121(83)$ $1.0(2.2)$ $0.10(27)$ $-0.7(15.9)$ $-0.1(28.6)$ $-0.7(9.1)$
$\Xi^{0}_{c}\to\Sigma^{0}\eta$ $0.020(14)$ $0.7(2.0)$ $0.02(12)$ $-0.03(2.31)$ $-0.98(69)$ $-0.2(3.8)$ $-0.04(17)$
$\Xi^{0}_{c}\to\Sigma^{0}\eta^{\prime}$ $0.0070(74)$ $1.0(2.6)$ $0.010(24)$ $-0.7(8.6)$ $-0.2(13.2)$ $-0.6(5.3)$ $0.05(51)$
$\Xi^{0}_{c}\to\Lambda\eta$ $0.015(40)$ $0.9(2.6)$ $0.02(15)$ $-0.8(1.6)$ $-0.5(1.1)$ $-0.3(3.8)$ $-0.12(59)$
$\Xi^{0}_{c}\to\Lambda\eta^{\prime}$ $0.004(56)$ $0.4(23.2)$ $0.05(23)$ $-0.4(7.8)$ $0.8(5.4)$ $-0.5(16.5)$ $-0.2(1.01)$
$\Xi^{0}_{c}\to n\eta$ $0.0008(34)$ $-0.2(1.7)$ $0.0025(90)$ $-0.6(2.3)$ $-0.38(65)$ $-0.7(2.2)$
$\Xi^{0}_{c}\to n\eta^{\prime}$ $0.0005(52)$ $0.8(9.4)$ $0.001(15)$ $-0.2(4.3)$ $-0.04(19.52)$ $-0.99(13)$
@@
$\Lambda^{+}_{c}\to\Sigma^{0}\pi^{+}$ $1.259(46)$ $-0.471(15)$ $1.261(49)$ $-0.472(15)$ $0.78(11)$ $-0.42(20)$
$\Lambda^{+}_{c}\to\Lambda\pi^{+}$ $1.305(47)$ $-0.7538(60)$ $1.290(49)$ $-0.7550(60)$ $-0.47(12)$ $-0.46(12)$
$\Lambda^{+}_{c}\to\Sigma^{+}\pi^{0}$ $1.273(46)$ $-0.469(15)$ $1.272(49)$ $-0.471(15)$ $0.77(11)$ $-0.42(20)$
$\Lambda^{+}_{c}\to pK_{S}^{0}$ $1.646(48)$ $0.41(12)$ $1.634(49)$ $0.12(26)$ $0.26(39)$ $-0.96(13)$
$\Lambda^{+}_{c}\to\Xi^{0}K^{+}$ $0.417(29)$ $0.960(17)$ $0.549(69)$ $0.02(15)$ $-0.011(77)$ $0.9997(43)$
$\Xi^{+}_{c}\to\Sigma^{+}K_{S}^{0}$ $0.79(21)$ $0.51(18)$ $0.043(11)$ $-0.96(21)$ $0.3(1.1)$ $-0.08(1.28)$ $0.01(29)$
$\Xi^{+}_{c}\to\Xi^{0}\pi^{+}$ $0.880(78)$ $-0.892(39)$ $1.90(68)$ $-0.82(10)$ $0.47(18)$ $0.31(15)$
$\Xi^{0}_{c}\to\Sigma^{0}K_{S}^{0}$ $0.056(16)$ $-0.55(26)$ $0.055(16)$ $-0.2(3.7)$ $-0.5(3.3)$ $0.8(2.7)$
$\Xi^{0}_{c}\to\Lambda K^{0}_{S}$ $0.248(29)$ $0.87(15)$ $0.331(59)$ $-0.41(59)$ $-0.75(51)$ $-0.51(39)$
$\Xi^{0}_{c}\to\Sigma^{+}K^{-}$ $0.189(38)$ $0.91(25)$ $0.182(40)$ $0.4(2.0)$ $-0.18(62)$ $-0.91(93)$
$\Xi^{0}_{c}\to\Xi^{-}\pi^{+}$ $1.17(18)$ $-0.701(45)$ $1.25(20)$ $-0.640(51)$ $0.16(77)$ $-0.75(16)$
$\Xi^{0}_{c}\to\Xi^{0}\pi^{0}$ $0.153(48)$ $-0.45(13)$ $0.62(13)$ $-0.82(24)$ $-0.49(69)$ $-0.30(56)$
$\Lambda^{+}_{c}\to\Sigma^{0}K^{+}$ $0.0376(18)$ $-0.9971(36)$ $0.0363(29)$ $-0.54(17)$ $-0.07(39)$ $-0.84(11)$ $-0.02(10)$
$\Lambda^{+}_{c}\to\Lambda K^{+}$ $0.0641(29)$ $-0.556(44)$ $0.0643(31)$ $-0.579(50)$ $0.809(40)$ $-0.10(17)$ $-0.02(14)$
$\Lambda^{+}_{c}\to\Sigma^{+}K^{0}_{S}/K^{0}_{L}$ $0.0311(29)$ $-0.64(10)$ $0.043(11)$ $-0.96(21)$ $0.3(1.1)$ $-0.08(1.28)$ $0.01(29)$
$\Lambda^{+}_{c}\to p\pi^{0}$ $0.0174(53)$ $-0.10(13)$ $0.0164(64)$ $0.3(2.5)$ $-0.5(2.5)$ $0.79(59)$
$\Lambda^{+}_{c}\to n\pi^{+}$ $0.0641(23)$ $0.547(42)$ $0.063(11)$ $-0.56(21)$ $-0.81(15)$ $0.18(27)$ $-0.33(28)$
$\Xi^{+}_{c}\to\Sigma^{0}\pi^{+}$ $0.3175(88)$ $-0.731(17)$ $0.303(19)$ $-0.676(58)$ $0.56(11)$ $-0.48(12)$ $-0.010(23)$
$\Xi^{+}_{c}\to\Lambda\pi^{+}$ $0.0211(32)$ $-0.19(18)$ $0.047(18)$ $0.60(23)$ $-0.47(43)$ $-0.64(31)$ $-0.23(26)$
$\Xi^{+}_{c}\to\Sigma^{+}\pi^{0}$ $0.273(18)$ $0.43(12)$ $0.256(87)$ $-0.13(35)$ $0.43(54)$ $-0.89(31)$ $0.073(93)$
$\Xi^{+}_{c}\to pK^{0}_{S}/K^{0}_{L}$ $0.164(15)$ $-0.451(82)$ $0.178(90)$ $-0.86(28)$ $0.2(1.1)$ $-0.5(1.0)$ $0.007(169)$
$\Xi^{+}_{c}\to\Xi^{0}K^{+}$ $0.1336(54)$ $0.382(33)$ $0.107(23)$ $-0.48(20)$ $-0.70(16)$ $0.53(20)$ $-0.39(32)$
$\Xi^{0}_{c}\to\Sigma^{0}\pi^{0}$ $0.00002(12)$ $-0.99(74)$ $0.028(37)$ $-0.88(64)$ $-0.38(70)$ $-0.3(1.2)$ $-0.13(14)$
$\Xi^{0}_{c}\to\Lambda\pi^{0}$ $0.0331(47)$ $0.74(11)$ $0.0129(65)$ $-0.84(33)$ $0.25(68)$ $-0.47(44)$ $-0.03(26)$
$\Xi^{0}_{c}\to\Sigma^{+}\pi^{-}$ $0.0118(25)$ $0.89(27)$ $0.0118(26)$ $0.4(1.9)$ $-0.17(59)$ $-0.92(82)$
$\Xi^{0}_{c}\to pK^{-}$ $0.0146(39)$ $0.79(32)$ $0.0157(45)$ $0.3(1.6)$ $-0.14(50)$ $-0.95(56)$
$\Xi^{0}_{c}\to\Sigma^{-}\pi^{+}$ $0.0630(97)$ $-0.769(44)$ $0.082(13)$ $-0.577(97)$ $0.02(72)$ $-0.817(85)$ $-0.016(38)$
$\Xi^{0}_{c}\to nK^{0}_{S}/K^{0}_{L}$ $0.0234(54)$ $0.40(25)$ $0.022(45)$ $0.20(48)$ $0.01(46)$ $-0.979(99)$ $-0.14(33)$
$\Xi^{0}_{c}\to\Xi^{-}K^{+}$ $0.0520(80)$ $-0.656(44)$ $0.0485(74)$ $-0.69(10)$ $0.03(86)$ $-0.73(12)$ $-0.024(56)$
$\Xi^{0}_{c}\to\Xi^{0}K^{0}_{S}/K^{0}_{L}$ $0.0103(20)$ $0.59(33)$ $0.009(20)$ $0.31(70)$ $0.02(70)$ $-0.95(23)$ $-0.28(56)$
$\Lambda^{+}_{c}\to pK^{0}_{L}$ $1.743(51)$ $0.48(12)$ $1.73(13)$ $0.18(30)$ $0.23(40)$ $-0.96(14)$
$\Lambda^{+}_{c}\to nK^{+}$ $0.001009(90)$ $-0.975(21)$ $0.00227(79)$ $-0.86(10)$ $0.50(18)$ $0.06(17)$
$\Xi^{+}_{c}\to\Sigma^{0}K^{+}$ $0.01142(30)$ $-0.9964(13)$ $0.01048(74)$ $-0.661(60)$ $-0.06(16)$ $-0.748(49)$
$\Xi^{+}_{c}\to\Lambda K^{+}$ $0.00441(18)$ $0.615(30)$ $0.00432(87)$ $-0.50(16)$ $0.813(80)$ $0.31(27)$
$\Xi^{+}_{c}\to\Sigma^{+}K^{0}_{L}$ $0.98(23)$ $0.74(15)$ $0.8(1.3)$ $0.8(1.7)$ $-0.5(2.9)$ $0.2(1.2)$
$\Xi^{+}_{c}\to p\pi^{0}$ $0.00067(25)$ $0.06(32)$ $0.00070(72)$ $0.6(1.6)$ $0.88(89)$ $-0.2(2.4)$
$\Xi^{+}_{c}\to n\pi^{+}$ $0.00607(39)$ $0.939(21)$ $0.00410(52)$ $0.04(29)$ $-0.02(15)$ $0.999(15)$
$\Xi^{0}_{c}\to\Sigma^{0}K_{L}^{0}$ $0.070(18)$ $-0.27(28)$ $0.079(55)$ $-0.4(2.5)$ $-0.6(2.6)$ $0.7(3.3)$
$\Xi^{0}_{c}\to\Lambda K^{0}_{L}$ $0.230(26)$ $0.91(12)$ $0.316(72)$ $-0.44(65)$ $-0.76(48)$ $-0.47(32)$
$\Xi^{0}_{c}\to p\pi^{-}$ $0.00087(24)$ $0.77(33)$ $0.00095(29)$ $0.3(1.5)$ $-0.14(49)$ $-0.95(51)$
$\Xi^{0}_{c}\to\Sigma^{-}K^{+}$ $0.00285(44)$ $-0.736(45)$ $0.00339(53)$ $-0.607(50)$ $0.15(73)$ $-0.78(15)$
$\Xi^{0}_{c}\to n\pi^{0}$ $0.00162(25)$ $0.993(32)$ $0.00043(58)$ $-0.9(1.2)$ $0.14(47)$ $0.4(2.9)$

### v3

@@ Acknowledgements.
$\Lambda^{+}_{c}\to p\pi^{0}$ $0.0156(75)$ $0.0174(53)$ $0.0160(71)$
$\Lambda^{+}_{c}\to pK_{S}^{0}$ $1.59(7)$ $0.2(5)$ $1.646(48)$ $0.41(12)$ $1.590(56)$ $0.05(27)$
$\Lambda^{+}_{c}\to pK_{L}^{0}$ $1.67(7)$ $1.646(48)$ $1.685(60)$
$\Lambda^{+}_{c}\to p\eta$ $0.158(11)$ $0.158(11)$ $0.154(11)$
$\Lambda^{+}_{c}\to p\eta^{\prime}$ $0.0484(91)$ $0.0464(61)$ $0.0513(75)$
$\Lambda^{+}_{c}\to\Lambda\pi^{+}$ $1.29(5)$ $-0.755(6)$ $1.305(47)$ $-0.7538(60)$ $1.292(48)$ $-0.7551(60)$
$\Lambda^{+}_{c}\to\Sigma^{0}\pi^{+}$ $1.27(6)$ $-0.466(18)$ $1.259(46)$ $-0.471(15)$ $1.259(49)$ $-0.472(15)$
$\Lambda^{+}_{c}\to\Sigma^{+}\pi^{0}$ $1.24(9)$ $-0.484(27)$ $1.273(46)$ $-0.469(15)$ $1.270(49)$ $-0.471(15)$
$\Lambda^{+}_{c}\to\Xi^{0}K^{+}$ $0.55(7)$ $0.01(16)$ $0.417(29)$ $0.545(69)$ 0.01(15)
$\Lambda^{+}_{c}\to\Lambda^{0}K^{+}$ $0.0642(31)$ $-0.58(5)$ $0.0641(29)$ $-0.556(44)$ $0.0646(30)$ $-0.578(50)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta$ $0.32(5)$ $-0.99(6)$ $0.327(49)$ $-0.9998(35)$ $0.328(49)$ $-0.978(57)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.41(8)$ $-0.46(7)$ $0.416(65)$ $-0.455(64)$ $0.379(71)$ $-0.446(66)$
$\Lambda^{+}_{c}\to\Sigma^{0}K^{+}$ $0.0370(31)$ $-0.54(20)$ $0.0376(18)$ $-0.9971(36)$ $0.0363(27)$ $-0.47(17)$
$\Lambda^{+}_{c}\to n\pi^{+}$ $0.066(13)$ $0.0641(23)$ $0.0646(91)$
$\Lambda^{+}_{c}\to\Sigma^{+}K_{S}^{0}$ $0.047(14)$ $0.0311(29)$ $0.0341(88)$
$\Xi^{+}_{c}\to\Xi^{0}\pi^{+}$ $1.6(8)$ $0.880(78)$ $1.03(30)$
$\Xi^{0}_{c}\to\Lambda K_{S}^{0}$ $0.32(6)$ $0.248(29)$ $0.342(58)$
$\Xi^{0}_{c}\to\Xi^{-}\pi^{+}$ $1.43(27)$ $-0.640(51)$ $1.17(18)$ $-0.701(45)$ $1.25(19)$ $-0.643(51)$
$\Xi^{0}_{c}\to\Xi^{-}K^{+}$ $0.039(11)$ $0.0520(80)$ $0.0488(75)$
@@
$\Lambda^{+}_{c}\to\Sigma^{+}\eta$ $0.327(49)$ $-0.9998(35)$ $0.332(77)$ $-0.95(21)$ $-0.3(1.7)$ $0.1(2.2)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.416(65)$ $-0.455(64)$ $0.409(99)$ $-0.409(88)$ $-0.70(59)$ $0.59(68)$
$\Lambda^{+}_{c}\to p\eta$ $0.154(10)$ $0.885(78)$ $0.162(21)$ $0.46(59)$ $-0.11(51)$ $-0.88(28)$ $-0.007(18)$
$\Lambda^{+}_{c}\to p\eta^{\prime}$ $0.0466(61)$ $-0.990(16)$ $0.056(15)$ $-0.90(34)$ $-0.43(68)$ $-0.04(49)$ $0.051(71)$
$\Xi^{+}_{c}\to\Sigma^{+}\eta$ $0.116(14)$ $0.91(11)$ $0.097(89)$ $0.50(45)$ $-0.04(1.71)$ $-0.87(24)$ $-0.020(57)$
$\Xi^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.117(16)$ $-0.418(71)$ $0.125(56)$ $-0.31(24)$ $-0.84(60)$ $0.4(1.2)$ $0.037(53)$
$\Xi^{+}_{c}\to p\eta$ $0.00908(67)$ $-0.054(61)$ $0.0105(41)$ $-0.12(13)$ $0.48(39)$ $-0.87(20)$
$\Xi^{+}_{c}\to p\eta^{\prime}$ $0.0091(10)$ $-0.9957(86)$ $0.0100(63)$ $-0.82(44)$ $-0.58(59)$ $-0.01(1.10)$
$\Xi^{0}_{c}\to\Xi^{0}\eta$ $0.177(75)$ $1.000(98)$ $0.107(99)$ $0.7(3.7)$ $-0.6(5.0)$ $-0.3(2.4)$
$\Xi^{0}_{c}\to\Xi^{0}\eta^{\prime}$ $0.121(83)$ $1.0(2.2)$ $0.10(21)$ $-0.7(2.9)$ $-0.2(16.8)$ $0.7(4.4)$
$\Xi^{0}_{c}\to\Sigma^{0}\eta$ $0.020(14)$ $0.7(2.0)$ $0.020(42)$ $-0.1(3.2)$ $-0.9(1.9)$ $0.4(3.7)$ $-0.026(61)$
$\Xi^{0}_{c}\to\Sigma^{0}\eta^{\prime}$ $0.0070(74)$ $1.0(2.6)$ $0.005(44)$ $-0.8(3.1)$ $-0.5(10.7)$ $0.4(11.5)$ $-0.04(28)$
$\Xi^{0}_{c}\to\Lambda\eta$ $0.015(40)$ $0.9(2.6)$ $0.011(16)$ $-0.3(7.4)$ $-1.0(2.1)$ $-0.06(1.33)$ $-0.020(63)$
$\Xi^{0}_{c}\to\Lambda\eta^{\prime}$ $0.004(56)$ $0.4(23.2)$ $0.01(20)$ $0.6(2.6)$ $-0.3(9.3)$ $-0.8(5.0)$ $-0.002(189)$
$\Xi^{0}_{c}\to n\eta$ $0.0008(34)$ $-0.2(1.7)$ $0.0017(33)$ $-0.2(1.8)$ $-0.5(1.4)$ $-0.82(73)$
$\Xi^{0}_{c}\to n\eta^{\prime}$ $0.0005(52)$ $0.8(9.4)$ $0.0010(32)$ $-0.8(5.4)$ $-0.4(11.2)$ $-0.4(2.7)$
@@
$\Lambda^{+}_{c}\to\Sigma^{0}\pi^{+}$ $1.259(46)$ $-0.471(15)$ $1.259(49)$ $-0.472(15)$ $0.75(13)$ $-0.46(20)$
$\Lambda^{+}_{c}\to\Lambda\pi^{+}$ $1.305(47)$ $-0.7538(60)$ $1.292(48)$ $-0.7551(60)$ $-0.480(75)$ $-0.446(81)$
$\Lambda^{+}_{c}\to\Sigma^{+}\pi^{0}$ $1.273(46)$ $-0.469(15)$ $1.270(49)$ $-0.471(15)$ $0.75(13)$ $-0.47(20)$
$\Lambda^{+}_{c}\to pK_{S}^{0}$ $1.646(48)$ $0.41(12)$ $1.590(56)$ $0.05(27)$ $0.19(64)$ $-0.98(13)$
$\Lambda^{+}_{c}\to\Xi^{0}K^{+}$ $0.417(29)$ $0.960(17)$ $0.545(69)$ $0.01(15)$ $-0.008(122)$ $0.9999(24)$
$\Xi^{+}_{c}\to\Sigma^{+}K_{S}^{0}$ $0.79(21)$ $0.51(18)$ $0.60(76)$ $0.7(1.6)$ $-0.5(2.5)$ $0.48(58)$
$\Xi^{+}_{c}\to\Xi^{0}\pi^{+}$ $0.880(78)$ $-0.892(39)$ $1.03(30)$ $-0.90(12)$ $0.43(24)$ $0.09(24)$
$\Xi^{0}_{c}\to\Sigma^{0}K_{S}^{0}$ $0.056(16)$ $-0.55(26)$ $0.055(16)$ $-0.03(2.90)$ $0.07(2.74)$ $0.997(97)$
$\Xi^{0}_{c}\to\Lambda K^{0}_{S}$ $0.248(29)$ $0.87(15)$ $0.342(58)$ $-0.27(73)$ $-0.81(87)$ $-0.5(1.1)$
$\Xi^{0}_{c}\to\Sigma^{+}K^{-}$ $0.189(38)$ $0.91(25)$ $0.185(40)$ $0.2(1.7)$ $-0.1(1.1)$ $-0.98(44)$
$\Xi^{0}_{c}\to\Xi^{-}\pi^{+}$ $1.17(18)$ $-0.701(45)$ $1.25(19)$ $-0.643(51)$ $0.06(1.39)$ $-0.76(12)$
$\Xi^{0}_{c}\to\Xi^{0}\pi^{0}$ $0.153(48)$ $-0.45(13)$ $0.59(12)$ $-0.78(19)$ $-0.4(1.3)$ $-0.50(76)$
$\Lambda^{+}_{c}\to\Sigma^{0}K^{+}$ $0.0376(18)$ $-0.9971(36)$ $0.0363(27)$ $-0.47(17)$ $-0.07(26)$ $-0.881(86)$ $0.003(23)$
$\Lambda^{+}_{c}\to\Lambda K^{+}$ $0.0641(29)$ $-0.556(44)$ $0.0646(30)$ $-0.578(50)$ $0.810(41)$ $-0.10(20)$ $0.014(20)$
$\Lambda^{+}_{c}\to\Sigma^{+}K^{0}_{S}/K^{0}_{L}$ $0.0311(29)$ $-0.64(10)$ $0.0341(88)$ $-0.76(42)$ $0.59(97)$ $-0.28(99)$ $0.063(91)$
$\Lambda^{+}_{c}\to p\pi^{0}$ $0.0174(53)$ $-0.10(13)$ $0.0160(71)$ $0.6(1.6)$ $-0.09(2.71)$ $0.79(95)$
$\Lambda^{+}_{c}\to n\pi^{+}$ $0.0641(23)$ $0.547(42)$ $0.0646(91)$ $-0.69(22)$ $-0.58(21)$ $0.43(15)$ $0.076(57)$
$\Xi^{+}_{c}\to\Sigma^{0}\pi^{+}$ $0.3175(88)$ $-0.731(17)$ $0.307(14)$ $-0.631(49)$ $0.578(97)$ $-0.52(14)$ $0.0(1.1)\times 10^{-17}$
$\Xi^{+}_{c}\to\Lambda\pi^{+}$ $0.0211(32)$ $-0.19(18)$ $0.044(14)$ $0.55(32)$ $-0.69(38)$ $-0.48(29)$ $-0.10(11)$
$\Xi^{+}_{c}\to\Sigma^{+}\pi^{0}$ $0.273(18)$ $0.43(12)$ $0.261(62)$ $0.05(24)$ $0.40(32)$ $-0.92(13)$ $0.009(18)$
$\Xi^{+}_{c}\to pK^{0}_{S}/K^{0}_{L}$ $0.164(15)$ $-0.451(82)$ $0.154(51)$ $-0.63(18)$ $0.49(97)$ $-0.60(68)$ $0.034(55)$
$\Xi^{+}_{c}\to\Xi^{0}K^{+}$ $0.1336(54)$ $0.382(33)$ $0.119(17)$ $-0.55(19)$ $-0.46(16)$ $0.696(92)$ $0.092(70)$
$\Xi^{0}_{c}\to\Sigma^{0}\pi^{0}$ $0.00002(12)$ $-0.99(74)$ $0.029(31)$ $-0.89(52)$ $-0.4(1.1)$ $-0.12(96)$ $-0.049(74)$
$\Xi^{0}_{c}\to\Lambda\pi^{0}$ $0.0331(47)$ $0.74(11)$ $0.015(12)$ $-0.83(50)$ $0.3(1.1)$ $-0.44(55)$ $-0.035(43)$
$\Xi^{0}_{c}\to\Sigma^{+}\pi^{-}$ $0.0118(25)$ $0.89(27)$ $0.0120(26)$ $0.2(1.6)$ $-0.1(1.0)$ $-0.98(39)$
$\Xi^{0}_{c}\to pK^{-}$ $0.0146(39)$ $0.79(32)$ $0.0162(37)$ $0.1(1.3)$ $-0.10(83)$ $-0.99(25)$
$\Xi^{0}_{c}\to\Sigma^{-}\pi^{+}$ $0.0630(97)$ $-0.769(44)$ $0.083(13)$ $-0.577(49)$ $0.06(1.25)$ $-0.815(97)$ $0.0(8.3)\times 10^{-18}$
$\Xi^{0}_{c}\to nK^{0}_{S}/K^{0}_{L}$ $0.0234(54)$ $0.40(25)$ $0.029(32)$ $0.18(49)$ $-0.54(81)$ $-0.82(58)$ $0.003(44)$
$\Xi^{0}_{c}\to\Xi^{-}K^{+}$ $0.0520(80)$ $-0.656(44)$ $0.0488(75)$ $-0.687(52)$ $0.07(1.48)$ $-0.72(14)$ $0.0(1.2)\times 10^{-17}$
$\Xi^{0}_{c}\to\Xi^{0}K^{0}_{S}/K^{0}_{L}$ $0.0103(20)$ $0.59(33)$ $0.013(15)$ $0.25(62)$ $-0.74(85)$ $-0.6(1.1)$ $0.006(93)$
$\Lambda^{+}_{c}\to pK^{0}_{L}$ $1.743(51)$ $0.48(12)$ $1.685(60)$ $0.11(30)$ $0.15(66)$ $-0.98(12)$
$\Lambda^{+}_{c}\to nK^{+}$ $0.001009(90)$ $-0.975(21)$ $0.00130(37)$ $-0.89(12)$ $0.43(24)$ $-0.17(23)$
$\Xi^{+}_{c}\to\Sigma^{0}K^{+}$ $0.01142(30)$ $-0.9964(13)$ $0.01104(52)$ $-0.699(58)$ $-0.07(13)$ $-0.712(52)$
$\Xi^{+}_{c}\to\Lambda K^{+}$ $0.00441(18)$ $0.615(30)$ $0.00374(52)$ $-0.34(19)$ $0.892(75)$ $0.30(28)$
$\Xi^{+}_{c}\to\Sigma^{+}K^{0}_{L}$ $0.98(23)$ $0.74(15)$ $0.77(85)$ $0.8(1.3)$ $-0.5(2.2)$ $0.17(67)$
$\Xi^{+}_{c}\to p\pi^{0}$ $0.00067(25)$ $0.06(32)$ $0.0008(15)$ $0.4(1.6)$ $0.7(1.2)$ $-0.6(2.2)$
$\Xi^{+}_{c}\to n\pi^{+}$ $0.00607(39)$ $0.939(21)$ $0.00407(51)$ $0.02(29)$ $-0.01(23)$ $0.9997(88)$
$\Xi^{0}_{c}\to\Sigma^{0}K_{L}^{0}$ $0.070(18)$ $-0.27(28)$ $0.074(47)$ $-0.3(2.3)$ $-0.1(2.7)$ $0.95(98)$
$\Xi^{0}_{c}\to\Lambda K^{0}_{L}$ $0.230(26)$ $0.91(12)$ $0.321(71)$ $-0.32(75)$ $-0.82(86)$ $-0.5(1.1)$
$\Xi^{0}_{c}\to p\pi^{-}$ $0.00087(24)$ $0.77(33)$ $0.00098(23)$ $0.1(1.3)$ $-0.09(79)$ $-0.99(23)$
$\Xi^{0}_{c}\to\Sigma^{-}K^{+}$ $0.00285(44)$ $-0.736(45)$ $0.00338(51)$ $-0.609(50)$ $0.06(1.32)$ $-0.79(11)$
$\Xi^{0}_{c}\to n\pi^{0}$ $0.00162(25)$ $0.993(32)$ $0.00045(29)$ $-0.7(1.1)$ $0.50(45)$ $0.5(1.9)$